# Autoencoding-Free Context Compression for LLMs via Contextual Semantic Anchors

### v3

@@ 4.2 Finetuning Results
Methods BioASQ DROP DouRC RACE RE TQA Avg
@@
4.3 Ablation Study
To verify the effectiveness of each key component and strategy in the SAC architecture, we conduct three groups of ablation studies, all performed under a $5\times$ compression ratio.
Component Ablation. As shown in Table 3, our ablation study clearly demonstrates the critical roles of the bidirectional attention and anchor embedding. Removing either component leads to substantial performance degradation in both in-domain (ID) and out-of-domain (OOD) settings. The bidirectional attention mechanism enables anchor tokens to more effectively integrate information from the entire context, producing compressed representations that are more beneficial for downstream tasks. Meanwhile, the anchor embedding provides explicit structural signals that guide the model to accurately identify and process these key tokens, thereby ensuring the effectiveness of information compression.
Table 3: Component ablation results. We report the average F1/EM performance of the model on in-domain (ID) and out-of-domain (OOD) tasks after removing the bidirectional attention (w/o mask) and the anchor embedding (w/o anchor). Full results on all tasks are provided in the Appendix B.2.
Methods ID OOD
TriviaQA HotpotQA Avg BioASQ TextbookQA Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Token Selection. As shown in Table 4, our ablation study investigates the effect of different token selection strategies on the performance of SAC. The results indicate that random selection (Random) significantly degrades performance, not only because the selected tokens lack importance, but also due to their positional randomness, which leads to insufficient global coverage and fails to effectively represent the context. In contrast, both information-based selection (Lingua-2) and our default strategy achieve near-optimal results, and both substantially outperform existing baselines in Tables 1 and 2. This demonstrates that the SAC architecture can effectively leverage and enhance any high-quality token selection strategy, rather than relying on a specific choice, highlighting the generality and robustness of the SAC framework.
Table 4: Token selection ablation results. This table demonstrates how different token selection strategies affect model performance, comparing Random selection, Lingua-2-based selection (Pan et al., 2024b), and our uniform selection (Zhao et al., 2025). We report average F1/EM scores across in-domain (ID) and out-of-domain (OOD) tasks, with comprehensive results for all individual tasks presented in the Appendix B.2.
Methods ID OOD
TriviaQA HotpotQA Avg BioASQ TextbookQA Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
SAC(Random) 59.24 50.22 58.84 41.86 59.04 46.04 43.18 30.59 45.36 29.27 44.27 29.93
SAC(Lingua-2) 64.55 55.13 67.05 49.74 65.80 52.44 44.49 31.91 51.46 32.07 47.98 31.99
AE Effect. As shown in Table 5, we compare the effect of introducing an autoencoding (AE) objective during training on the performance of SAC. Traditional context compression methods employ AE tasks to force independent compression tokens to attend to the original context for reconstruction. However, we find that the AE objective itself has inherent limitations, as its reconstruction target is misaligned with downstream tasks. The experimental results validate this observation: training with only the AE objective leads to a substantial performance drop, and even when combined with the LM objective, the performance still lags behind the full SAC model. This highlights the architectural advantage of SAC: since anchor tokens are naturally semantically aligned with the original context, our method does not require AE objectives to force learning. Instead, SAC effectively aggregates contextual information into anchor token representations solely through anchor embeddings and bidirectional attention. It is worth noting that the ablation experiments in ICAE demonstrate that combining autoencoding tasks with language modeling tasks yields better results (Ge et al., 2024). However, our reproduction on 500xCompressor does not fully support this finding. Specifically, under 15x and 51x compression ratios, 500xCompressor achieves better in-distribution (ID) results when using language modeling tasks alone, with other scenarios being exceptions. This observation raises questions regarding the necessity of autoencoding tasks and suggests that autoencoding may not be entirely essential for context compression methods.
Methods ID OOD
5x 15x 51x 5x 15x 51x
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
5 ANALYSIS
5.1 Attention Visualization
To understand the unique behavior of compressed models, we analyzed the attention patterns of the final layer at a 5x compression rate. Attention maps for other compression rates can be found in Appendix  C.2.
As observed in Figure 4, The attention map for the 500xCompressor exhibits a distinct anti-diagonal trend. To complete the autoencoding task, the model must condense the entire original sequence into these remaining compressed tokens. This forces later compressed tokens to break locality constraints and actively seek out and attend to distant but important tokens in the sequence. In contrast, the SAC model demonstrates a clear diagonal pattern, where its anchor tokens effectively attend to their neighboring original context tokens, showing a better ability to focus on local information.
Figure 4: Attention maps of different models finetuned under a 5x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression/anchor tokens.
5.2 Representation Analysis
Key Representation Analysis.In the Key representation space (see Figure 5), the compression tokens (orange) of SAC and EPL are distributed relatively close to the context tokens (blue), while the compression tokens of 500xCompressor are clearly separated from the context tokens. This discrepancy arises from the architectural design of each method. Specifically, although the positional IDs of 500xCompressor’s additional compression tokens are contiguous, their semantics are not aligned, leading to a complete separation in the Key space. In contrast, EPL modifies the positional IDs of its additional compression tokens to share the same rotational angle (RoPE) as the original context tokens, thereby reducing the distance between them. However, in SAC, the anchor tokens are directly embedded within the original context, and their representations maintain close semantic ties with the context tokens from the outset, naturally preventing significant representational divergence.
Value Representation Analysis. In the Value representation space (see Figure 5), the anchor tokens of SAC are uniformly distributed across all regions with the Value representations of the context tokens, without forming independent sub-clusters. This suggests that SAC’s anchor embedding strategy allows for compressed Value representations that more closely match the distribution of the original Value space. In contrast, although EPL’s compression tokens also overlap with the context tokens, their distribution is less complete than SAC’s: they appear relatively sparse in the core regions and show a slight clustering tendency at the boundaries. This indicates that EPL’s Value representations still exhibit a degree of semantic shift relative to the original Value space, which is even more pronounced in the 500xCompressor.
6 Conclusion
This paper proposes a novel, autoencoding-free context compression method, Semantic-Anchor Compression (SAC), designed to address the performance degradation in downstream tasks caused by context-agnostic compression tokens and autoencoding objectives in existing context compression methods. Unlike traditional context compression approaches, SAC does not rely on training compression tokens to reconstruct the original input. Instead, it directly selects representative anchor tokens from the context and aggregates contextual information into their key-value (KV) representations via a bidirectional attention mechanism. This approach effectively compresses lengthy contexts while avoiding any impairment to the language model’s original language modeling capabilities. Experiments on multiple question answering tasks demonstrate that SAC achieves a high compression ratio and significantly outperforms existing compression methods, highlighting its superiority in balancing compression efficiency and model performance.
7 Reproducibility statement
We declare that the work presented in this paper is reproducible. We provide a link to our anonymous source code as supplementary material: https://anonymous.4open.science/r/SAC-E32C. This code can be used to reproduce the experimental results. The repository includes detailed instructions for environment setup, running experiments, data processing, and result evaluation.
Appendix A Experiment Details
We perform pretraining and finetuning using bf16 precision on 8 NVIDIA RTX 3090 GPUs (24GB). For pretraining, we randomly sample data from the SlimPajama-6B dataset with a token length ranging from 510 to 2040. This data is then split into two halves: one for the auto-encoding (AE) task and the other for the language modeling (LM) task (the AE half is discarded for models without the AE objective). For downstream tasks, we process the MRQA dataset into a (Context, Question, Answer) format for finetuning. Detailed hyperparameters can be found in Table 6.
Table 6: Hyperparameters for training
Hyperparameter Value
Optimizer AdamW
Betas (0.9, 0.95)
Weight decay 0.1
Learning rate 1e-4 (pretrain)
5e-5 (finetuning)
Scheduler Constant
Batch size 16
Warmup 300
Training steps 20k (pretrain)
20k (finetuning)
Clip norm 2.0
Appendix B Detailed Results
B.1 Pretraining Results
Methods LM-PPL Training Time(h)
ICAE 12.35 3.85
500xCompress 11.83 3.60
EPL 10.88 3.60
SAC 10.79 2.66
B.2 Ablation Results
In the main text, we have discussed the significant performance gains of SAC over all baseline methods. To provide more detailed evidence, we present the full ablation study results here. As shown in Table 8 and Table 9, our conclusion holds not only in terms of average performance but is also consistently validated on each individual dataset.
Table 8: Ablation studies for SAC under a 5x compression rate on the in-domain dataset are conducted in three sets: component ablation, token selection, and the influence of the auto-encoding (AE) task.
Methods SQuAD NewsQA TriviaQA SearchQA HotpotQA NQ Avg
@@ B.2 Ablation Results
Table 9: Ablation studies for SAC under a 5x compression rate on the out-of-domain dataset are conducted in three sets: component ablation, token selection, and the influence of the auto-encoding (AE) task.
Methods BioASQ DROP DouRC RACE RE TQA Avg
@@
Appendix C Visualization Analysis
C.1 Training Curves Analysis
Figure 6: Supervised finetuning loss curves. The figure illustrates the training loss trajectories of different models under three compression ratios: 5x, 15x, and 51x.
C.2 Attention Analysis
At a lower 5x compression rate, as shown in Figure 7, the attention map of EPL presents a clear positive diagonal, indicating that its compressed tokens primarily attend to local tokens. In contrast, the attention map of 500xCompressor appears more diffused, while our SAC model exhibits a sparse and highly focused attention pattern, with its anchor tokens attending to only a few key original context tokens. This phenomenon becomes more pronounced with increasing compression rates, being most evident at the 51x compression rate in Figure 9, which strongly demonstrates the robustness of SAC in extreme compression environments.
Figure 7: Attention maps of different models finetuned under a 5x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression tokens.
Figure 8: Attention maps of different models finetuned under a 15x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression/anchor tokens.
Figure 9: Attention maps of different models finetuned under a 51x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression tokens.
Appendix D The Use of Large Language Models

### v4

@@ Appendix D The Use of Large Language Models
Methods BioASQ DROP DouRC RACE RE TQA Average
@@
[…]
Methods SQuAD NewsQA TriviaQA SearchQA HotpotQA NQ Average
@@
[…]
Methods BioASQ DROP DouRC RACE RE TQA Average
@@
Table 18: Experimental Results of the 1B/3B/8B Models on the In-Domain QA Tasks.
Methods SQuAD NewsQA TriviaQA SearchQA HotpotQA NQ Average
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
EPL(1B) 44.58 27.91 33.34 16.69 56.16 47.09 66.36 54.13 54.88 38.38 53.80 35.71 51.52 36.65
SAC(1B) 47.43 30.25 36.55 18.07 61.13 52.19 68.97 56.76 58.83 41.86 56.79 38.88 54.95 39.67
EPL(3B) 62.43 42.65 46.72 26.14 68.83 59.92 73.61 61.56 68.50 51.03 63.59 44.43 63.95 47.62
SAC(3B) 63.11 43.51 49.23 28.80 69.65 60.85 73.88 62.29 68.87 51.40 65.98 47.18 65.12 49.01
EPL(8B) 65.79 45.55 51.18 28.96 72.40 63.49 74.86 63.23 70.98 53.65 67.24 48.71 67.08 50.60
SAC(8B) 64.92 44.80 49.77 29.20 72.37 63.70 76.92 65.29 70.80 54.25 67.77 49.21 67.09 51.08
Table 19: Experimental Results of the 1B/3B/8B Models on the Out-of-Domain QA Tasks.
Methods BioASQ DROP DouRC RACE RE TQA Average
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
EPL(1B) 40.52 28.52 32.16 22.29 25.70 16.39 20.97 4.01 59.75 46.34 41.31 25.42 36.74 23.83
SAC(1B) 41.31 28.66 36.72 27.48 28.94 18.99 23.35 4.90 61.04 47.90 44.21 28.21 39.26 26.02
EPL(3B) 46.76 32.98 46.20 36.46 36.81 25.32 33.80 7.57 65.73 53.70 55.46 34.86 47.46 31.82
SAC(3B) 47.96 34.51 48.46 38.39 42.21 29.85 33.61 7.72 75.11 62.89 55.53 35.00 50.48 34.73
EPL(8B) 52.52 37.57 49.98 39.92 40.09 27.24 35.22 7.27 69.77 57.84 57.36 36.66 50.82 34.42
SAC(8B) 48.54 34.04 51.55 40.65 41.44 29.18 35.52 7.72 78.91 67.71 57.87 36.26 52.31 35.93
Table 20: Model Performance on 32K Long-Context Summarization Tasks.
Methods QMSum GovReport Average
SAC 14.95 22.03 18.49
SAC(ae+lm) 15.32 20.15 17.74
EPL 14.82 20.40 17.61
Table 21: Model Performance on 24K Long-Context QA Tasks.
Methods MultiFieldQA NarrativeQA Qasper Sigle-Doc-Avg. 2WikiMQA MuSiQue HotpotQA Multi-Doc-Avg.
SAC 17.75 11.33 11.67 13.58 24.78 12.36 26.03 21.06
SAC(ae+lm) 17.38 10.49 9.05 12.31 22.47 7.76 23.88 18.04
EPL 14.34 10.50 7.58 10.81 21.99 8.43 23.72 18.05
Figure 6: Attention maps of different models finetuned under a 5× compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression tokens.
Figure 7: Attention maps of different models finetuned under a 15× compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression/anchor tokens.
Figure 8: Attention maps of different models finetuned under a 51× compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression tokens.
Figure 9: Gradient Cosine Similarity between AutoEncoding (AE) Loss and Language Modeling (LM) Loss.
Figure 10: Efficiency and Performance Trade-off Curves on In-Domain (ID) Tasks.